Josh is 4 times as old as Nick. in 4 years the sum of their ages will equal 18. how old is nick now?

Answer by addingup(3677)

J = 4N
J+4+N+4 = 18 substitute for J:
4N+4+N+4 = 18
5N + 8 = 18
5N = 10
N = 2
Nick is 2 years old
Josh is 2*4 = 8 years old